# Smooth Deep Saliency

### v1

@@ Abstract
In this work, we investigate methods to reduce the noise in deep saliency maps coming from convolutional downsampling, with the purpose of explaining how a deep learning model detects tumors in scanned histological tissue samples. Those methods make the investigated models more interpretable for gradient-based saliency maps, computed in hidden layers. We test our approach on different models trained for image classification on ImageNet1K, and models trained for tumor detection on Camelyon16 and in-house real-world digital pathology scans of stained tissue samples. Our results show that the checkerboard noise in the gradient gets reduced, resulting in smoother and therefore easier to interpret saliency maps.
Machine Learning, Explainable AI, Digital Pathology, Computer Vision
@@ 1 Introduction
For critical applications, such as in the healthcare sector, it is important to gain an insight into how the model arrives at its decision. To gain such an understanding, saliency methods can be used, ideally highlighting the structures on which the model relies to support its prediction. Saliency methods are usually applied at either the input or the last convolutional layer (Selvaraju et al., 2017). Their application in the input layer leads to noisy results, and in the final convolution layer, the saliency map lacks detail. In this paper, we focus on the middle ground and apply saliency methods in hidden layers. This gives us smoother results while preserving the details. As a result, we can use high-quality saliency maps to identify individual cell nuclei, which are used by trained models to detect tumors. On the one hand, using hidden layers also allows us to observe the transition from lower-level features, like the outline of cell nuclei in a Resnet18 (He et al., 2016) trained on Camelyon16 (Ehteshami Bejnordi et al., 2017) data or individual cell nuclei for a segmentation model trained on digital pathology (Digipath) data at a lower magnification, highlighted in earlier layers. On the other hand, it also gives access to higher-level features in deeper layers, e.g. segmenting the whole cell nuclei for the Resnet18 model or structures formed by cells for the model trained on Digipath data.
@@
A problem in computing gradient-based saliency maps in hidden layers of a model that uses convolutional downsampling is the checkerboard noise this downsampling introduces to its gradient, similar to what transposed convolutions exhibit in the forward pass (Odena et al., 2016). This noise dominates the resulting saliency map and makes them very difficult to interpret. Therefore, in this paper, we investigate several methods to remove that checkerboard noise, which allows us to use saliency methods in hidden layers and test them on several models and datasets.
@@
With a stride 2 convolution (downsampling), not every input pixel gets multiplied by every weight. Due to the stride of 2, every other pixel is skipped, which can introduce a checkerboard pattern with a periodicity of 2 in the gradient (see Section 2.1). Since that noise is coming from the convolutional downsampling, we can mitigate it by changing the backward pass of each convolutional downsampling layer, in order to let each input pixel receive a gradient by each weight. The other method we use is to train a bilinear surrogate model, where we replace each convolutional downsampling with a bilinear downsampling combined with two trainable stride 1 convolutions, one before and after the bilinear downsampling. These are then trained to match the behavior of the original model. That way, we can compare the accuracy and predictions of the surrogate model to the original model (by only changing the backward pass of the convolutional layer, but we no longer know whether we still explain the original model). The output of each surrogate path, consisting of the two stride 1 convolutions with the bilinear downsampling in between, should match the output of the convolutional downsampling it replaces concerning the L1 loss. We train this bilinear surrogate model on the training data of the original model. We show that the accuracy and predictions of the surrogate model closely match those of the original model, while the generated saliency maps have less noise, as measured by total variation.
@@ 2 Related Work
Saliency maps are used to highlight structures in the input that the model utilized for its prediction. The saliency methods could be divided into three groups. First, those that only utilize activations like CAM (Oquab et al., 2015), secondly those that also utilize gradients of backpropagation like Grad-CAM (Selvaraju et al., 2017), DeepLift (Shrikumar et al., 2019), Integrated Gradients (Sundararajan et al., 2017), or simply using the gradient (Simonyan et al., 2014). And lastly pertubation base methods like RISE (Petsiuk et al., 2018), which are usually applied at the input. Pertubation-based methods generally require several evaluations of the model, which makes them slow. To get gradient-based methods more stable SmoothGrad (Smilkov et al., 2017) can be employed, which runs the saliency method several times each time adding Gaussian noise to the input and then returning the average saliency map.
Gradient-based saliency methods are usually applied at either the input layer or the last convolutional layer, like (Selvaraju et al., 2017), whereas we focus on computing the saliency maps from hidden layers of the model. (Olah et al., 2018) also computes saliency maps at hidden layers and couples them with feature visualization. To mitigate noise in the saliency maps, they employ Gaussian blurring. (Rao et al., 2022) also used Gaussian blurring to smooth saliency maps of hidden layers. We limit ourselves to models that utilize convolutional downsampling and, instead of using Gaussian blurring, we work on removing the source of the noise.
(Olah et al., 2017) visualized neurons of a CNN, and observed that the visualizations become more high-level the deeper the neuron is located in the model. We observe something similar in the attribution for the Camelyon16 and Digipath models, where in earlier layers the cell nuclei are outlined and in deeper layers segmented (Camelyon16). And the saliency map for the Digipath model highlights individual cell nuclei in earlier layers and structures formed by them in deeper layers.
@@ 2.1 Checkerboard Noise from Stride 2 Convolution
With a stride 2 convolution, not every pixel is multiplied by every weight. Due to the stride of 2, every other pixel is skipped. As an example, let $x$ be a matrix of ones with a size of $16\times 16$, $k=\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ and * the operator for a stride 2 convolution. Figure 1 shows the result of calculating $\nabla_{x}(\sum x*k)$, i.e. the gradient towards $x$ when computing the sum of the convolution of $x$ with kernel $k$. We can see a periodic checkerboard pattern in the gradient. The black pixels were multiplied by $-1.0$ from the kernel and, therefore, received a gradient of -1.0. In contrast, the white pixels were multiplied by the $1.0$ value from the kernel and received a gradient of 1.0. That is the noise pattern we mitigate with our methods.
@@ 3.1 Overview
  • Bilinear Surrogate: Replacing each convolutional downsampling with a bilinear surrogate path (consisting of two stride 1 convolutions with a bilinear downsampling in between them), as shown in Figure 2. This approach requires training data, in order to train the surrogate paths.
@@
In training of the bilinear surrogate model, we replace each convolutional downsampling individually (i.e. we only use the convolutional downsampling in the forward pass and not the bilinear surrogate one). But in the evaluation, we use them all together, as visible from Figure 3. We use the L1 loss as a loss function in training.
@@ 3.2 Evaluation
With our bilinear surrogate method, we aim to achieve two goals:
@@
To validate whether we succeeded in reducing the checkerboard noise in the gradient, we can visually inspect some saliency maps. Further, we also run total variation (TV) on the saliency maps over a validation dataset. To evaluate how well we match the prediction behavior of the original model, we compare the accuracy and/or difference in prediction for only the target class and all classes of the original and the surrogate model on a validation dataset. For the difference in prediction, we use the output of the model after softmax. Then we compute the difference in prediction twice: Once only for the target class and once for all classes. For only the target class that would be: Let $\tilde{y}$ the prediction of the original model and $\hat{y}$ be the prediction of the bilinear surrogate model and $c$ be the target class, then the difference in prediction would be: $|\tilde{y}_{c}-\hat{y}_{c}|$. Whereas for all classes we would do: $\sum_{c}|\tilde{y}_{c}-\hat{y}_{c}|$, where we compute the sum over all classes.
For the segmentation model on the Digipath data, we often have more than one class per image. Therefore, we compute the difference in prediction for the target class differently. For each sample, we only consider the non-background label with the majority of pixels. If the whole target is background, we skip that sample. Then, from the target and the original prediction, we build a binary mask. In that binary mask, we set a pixel to 1 if the original prediction (which is after softmax) for the target label at that pixel is larger than 0.1. Otherwise, we set the pixel value to 0. Finally, we report the weighted L1 loss between the prediction of the original model and the surrogate model (respectively the prediction of the forward hooked model), based on the mask, i.e. let M be the binary mask, then we do $\left(\sum_{c}|\tilde{y}_{c}-\hat{y}_{c}|\right)\cdot\frac{1}{\sum M}$
@@ 3.3 Computing the Saliency Map
We use two methods to compute the saliency maps. One is DeepLift, with a black image (zeros) as the baseline. The other is simply computing the gradient. We combine both with SmoothGrad, where we use 20 samples and Gaussian noise with a standard deviation of $\sigma=0.2$ as values for the parameters. The saliency map is computed using the logits output of the model (before softmax or sigmoid). Let $f$ be a function (in our case either DeepLift or gradient) that takes the image and baseline as input and returns the saliency map, $x$ the image and $b$ the baseline and $z^{(i)}\sim\mathcal{N}(0,0.2)$. With SmoothGrad we compute the mean $\frac{1}{n}\sum_{i=1}^{n}f(x+z^{(i)})$, with $n=20$ in our case.
@@ 3.4 Total Variation
$\epsilon$ is set to $10^{-6}$ and is used to avoid division by zero (i.e. if $z_{c}^{\prime}$ is zero everywhere). Equation 2 is the resulting total variation reported in the experiments section.
@@ 4.1.1 Hardware and Software
For all our experiments, we use a Linux server with 8 Nvidia RTX 2080 Ti GPUs. On the software side, we use PyTorch (Paszke et al., 2019) version 1.13.1 with CUDA version 11.6 and torchvision version 0.14.1. For computing the saliency maps we use the Captum library (Kokhlikyan et al., 2020) version 0.6.0.
@@ 4.1.2 Training
In this paragraph we describe the training setup we use in order to train the bilinear surrogate model. The forward and backward hook approaches work on the original model, and require no additional training. All networks in our experiments are running in eval mode. We only train the bilinear surrogate paths, each consisting of two 3x3 stride 1 convolutions, with a bilinear downsampling in between them. The rest of the network is kept frozen. On the ImageNet1K (Deng et al., 2009) and Digipath data we train the bilinear surrogate for 10 epochs, and on the Camelyon16 data we train it for 2 epochs, as shown in Table 1. As optimizer we use Adam (Kingma & Ba, 2015) with a learning rate of 0.001. As batch size we use 64 for training the surrogate models on the ImageNet1K data, 512 for the Camelyon16 data (due to class imbalance between tumor and no tumor we use a larger batch size here) and 16 for the Digipath data (due to the larger patch size of 1536 we use a smaller batch size here).
@@
Training Samples 1 281 16712811671\,281\,1671 281 167 99 410 6159941061599\,410\,61599 410 615 129 634129634129\,634129 634
Validation Samples 50 0005000050\,00050 000 50 0005000050\,00050 000 $899$
@@ 4.1.3 Datasets
  • ImageNet1K. We evaluate the accuracy and the difference in prediction between the original and surrogate model on the 50 0005000050\,00050 000 validation images. The total variation is evaluated on every hundredth image of the validation data, resulting in 500 images.
  • Camelyon16. The difference in prediction is evaluated on all the 46 267 8994626789946\,267\,89946 267 899 validation samples. The total variation is evaluated on 500 samples randomly chosen out of the validation samples.
@@ 4.1.4 Models
On the ImageNet1K data we use a ResNet34 and a ResNet50 both openly available from the pytorch model zoo. Further, we use an adversarially robust ResNet50 from (Santurkar et al., 2019). For the Camelyon16 data we use an openly available ResNet18 from the monai model zoo. And for the Digipath data we use an in-house U-Net segmentation model with a ResNet34 backbone.
4.2 Results
In this section we describe the results of our experiments. First we show the accuracy of the original model, the surrogate model and the forward hook approach, as well as the difference in prediction compared to the original model.
4.2.1 Accuracy and Prediction Difference
@@
Table 2 shows the accuracy of the original, bilinear surrogate and forward hooked models, on the ImageNet1K data. For the bilinear surrogate models we observe a maximum decrease of accuracy of 0.17 for the ResNet34 model, and a minimum decrease of 0.03 for the robust ResNet50 model. The forward hooked models have a larger decrease of accuracy, especially in the case of the ResNet34 model where the accuracy decreases by 20.04.
@@
In Table 3 we evaluate how the prediction of the models changes, once for the sum over all predicted classes and once only for the target class. We report the differences in softmax values times 100 (i.e. a reported difference of 9.52 for ResNet34 would mean a difference of 0.0952). The first three models (ResNet34, ResNet50 and ResNet50 Robust) are originally trained on ImageNet1K, as are the surrogate models. Consequently, we evaluate the difference in prediction for them on the ImageNet1K validation data. The ResNet18 Camelyon16 model and surrogate model are trained on Camelyon16, therefore we evaluate on the validation data of Camelyon16. And the ResNet34 Digipath model and surrogate model are trained on the in-house Digipath dataset, and evaluated on data not seen in training. For the bilinear surrogate models we observe the maximum change of prediction for the sum of all classes and only for the target class of 9.52 respectively 2.39 for the Resnet34 model. That aligns with the drop of accuracy being largest for the ResNet34 surrogate model. We observe the smallest change in prediction of 0.40 for the ResNet18 Camelyon16 model. Since we only have a single class there, the result for ’all classes’ and ’only target class’ is the same. In the case of the forward hooked models, the predictions differ more strongly. Here the largest change in prediction of 95.84 for the sum over all classes and 30.26 for only the target class is again in the ResNet34 model. The smallest change is 3.63 also again for the ResNet18 Camelyon16 model. Overall the bilinear surrogates matche the accuracy and prediction of the original models much closer compared to the forward hooked models.
Table 4: Total variation of the saliency map
Network Layer DeepLift Gradient
Original Bilinear Surrogate Backward Hook Forward Hook Original Bilinear Surrogate Backward Hook Forward Hook
ResNet34 Layer1 2.711 $\pm$ 0.059 2.110 $\pm$ 0.096 $1.858\pm 0.114$ $\textbf{1.824}\pm 0.117$ 2.627 $\pm$ 0.025 1.846 $\pm$ 0.052 $\textbf{1.301}\pm 0.047$ $1.335\pm 0.044$
Layer2 2.132 $\pm$ 0.066 1.860 $\pm$ 0.089 $1.751\pm 0.098$ $\textbf{1.678}\pm 0.085$ 1.713 $\pm$ 0.039 1.285 $\pm$ 0.058 $0.990\pm 0.047$ $\textbf{0.968}\pm 0.041$
Layer3 2.165 $\pm$ 0.049 1.879 $\pm$ 0.062 $1.775\pm 0.071$ $\textbf{1.693}\pm 0.061$ 1.894 $\pm$ 0.051 1.326 $\pm$ 0.048 $1.054\pm 0.038$ $\textbf{1.025}\pm 0.036$
ResNet50 Layer1 2.774 $\pm$ 0.055 2.156 $\pm$ 0.090 $1.777\pm 0.110$ $\textbf{1.757}\pm 0.105$ 2.660 $\pm$ 0.030 1.986 $\pm$ 0.060 $1.439\pm 0.065$ $\textbf{1.425}\pm 0.061$
Layer2 2.529 $\pm$ 0.066 2.147 $\pm$ 0.100 $1.980\pm 0.111$ $\textbf{1.881}\pm 0.100$ 2.318 $\pm$ 0.043 1.860 $\pm$ 0.082 $1.584\pm 0.090$ $\textbf{1.561}\pm 0.083$
Layer3 2.400 $\pm$ 0.051 2.171 $\pm$ 0.051 $2.064\pm 0.075$ $\textbf{2.000}\pm 0.066$ 2.315 $\pm$ 0.060 1.894 $\pm$ 0.076 $1.720\pm 0.082$ $\textbf{1.664}\pm 0.074$
ResNet50 Robust Layer1 2.044 $\pm$ 0.099 1.702 $\pm$ 0.117 $1.539\pm 0.155$ $\textbf{1.493}\pm 0.149$ 1.934 $\pm$ 0.051 1.397 $\pm$ 0.076 $1.108\pm 0.084$ $\textbf{1.095}\pm 0.077$
Layer2 2.209 $\pm$ 0.077 1.815 $\pm$ 0.105 $1.678\pm 0.115$ $\textbf{1.600}\pm 0.104$ 2.095 $\pm$ 0.045 1.502 $\pm$ 0.071 $1.252\pm 0.072$ $\textbf{1.218}\pm 0.065$
Layer3 2.360 $\pm$ 0.046 1.986 $\pm$ 0.070 $1.824\pm 0.078$ $\textbf{1.776}\pm 0.069$ 2.347 $\pm$ 0.067 1.713 $\pm$ 0.073 $1.507\pm 0.074$ $\textbf{1.468}\pm 0.067$
ResNet18 Camelyon16 Layer1 2.511 $\pm$ 0.041 1.917 $\pm$ 0.079 $1.718\pm 0.081$ $\textbf{1.671}\pm 0.076$ $2.254\pm 0.038$ $1.531\pm 0.061$ $1.146\pm 0.060$ $\textbf{1.105}\pm 0.054$
Layer2 2.328 $\pm$ 0.054 1.884 $\pm$ 0.081 $1.723\pm 0.076$ $\textbf{1.671}\pm 0.074$ $1.988\pm 0.060$ $1.361\pm 0.056$ $1.030\pm 0.050$ $\textbf{1.005}\pm 0.047$
Layer3 2.170 $\pm$ 0.050 1.620 $\pm$ 0.050 $1.516\pm 0.046$ $\textbf{1.433}\pm 0.054$ $2.189\pm 0.063$ $1.296\pm 0.042$ $1.048\pm 0.034$ $\textbf{1.001}\pm 0.034$
ResNet34 Digipath Layer1 2.864 $\pm$ 0.023 1.965 $\pm$ 0.120 $1.644\pm 0.179$ $\textbf{1.640}\pm 0.195$ $2.713\pm 0.031$ $1.741\pm 0.042$ $1.177\pm 0.045$ $\textbf{1.120}\pm 0.044$
Layer2 2.642 $\pm$ 0.040 1.811 $\pm$ 0.212 $1.601\pm 0.210$ $\textbf{1.508}\pm 0.205$ $2.391\pm 0.062$ $1.357\pm 0.119$ $0.902\pm 0.096$ $\textbf{0.880}\pm 0.100$
Layer3 2.395 $\pm$ 0.055 1.249 $\pm$ 0.136 $1.080\pm 0.129$ $\textbf{1.031}\pm 0.114$ $2.444\pm 0.060$ $1.026\pm 0.092$ $0.704\pm 0.069$ $\textbf{0.698}\pm 0.059$

### v3

@@ Abstract
In this work, we investigate methods to reduce the noise in deep saliency maps coming from convolutional downsampling. Those methods make the investigated models more interpretable for gradient-based saliency maps, computed in hidden layers. We evaluate the faithfulness of those methods using insertion and deletion metrics, finding that saliency maps computed in hidden layers perform better compared to both the input layer and GradCAM. We test our approach on different models trained for image classification on ImageNet1K, and models trained for tumor detection on Camelyon16 and in-house real-world digital pathology scans of stained tissue samples. Our results show that the checkerboard noise in the gradient gets reduced, resulting in smoother and therefore easier to interpret saliency maps. [^1] The code is available at https://github.com/rherdt185/smooth-deep-saliency
Keywords:
XAI Computer Vision Deep Learning
@@ 1 Introduction
For critical applications, such as in the healthcare sector, it is important to gain an insight into how the model arrives at its decision. To gain such an understanding, saliency methods can be used, ideally highlighting the structures on which the model relies to support its prediction. Saliency methods are usually applied at either the input or the last convolutional layer [16]. Their application in the input layer leads to noisy results, and in the final convolution layer, the saliency map lacks detail. In this paper, we focus on the middle ground and apply saliency methods in hidden layers. This gives us smoother results while preserving the details. As a result, we can use high-quality saliency maps to identify individual cell nuclei, which are used by trained models to detect tumors. On the one hand, using hidden layers also allows us to observe the transition from lower-level features, like the outline of cell nuclei in a Resnet18 [4] trained on Camelyon16 [3] data or individual cell nuclei for a segmentation model trained on digital pathology (Digipath) data at a lower magnification, highlighted in earlier layers. On the other hand, it also gives access to higher-level features in deeper layers, e.g. segmenting the whole cell nuclei for the Resnet18 model or structures formed by cells for the model trained on Digipath data.
@@
A problem in computing gradient-based saliency maps in hidden layers of a model that uses convolutional downsampling is the checkerboard noise this downsampling introduces to its gradient, similar to what transposed convolutions exhibit in the forward pass [7]. This noise dominates the resulting saliency map and makes them very difficult to interpret. Therefore, in this paper, we investigate several methods to remove that checkerboard noise, which allows us to use saliency methods in hidden layers and test them on several models and datasets.
@@
With a stride 2 convolution (downsampling), not every input pixel gets multiplied by every weight. Due to the stride of 2, every other pixel is skipped, which can introduce a checkerboard pattern with a periodicity of 2 in the gradient (see Sec. 2.1). Since that noise is coming from the convolutional downsampling, we can mitigate it by changing the backward pass of each convolutional downsampling layer, in order to let each input pixel receive a gradient by each weight. The other method we use is to train a bilinear surrogate model, where we replace each convolutional downsampling with a bilinear downsampling combined with two trainable stride 1 convolutions, one before and after the bilinear downsampling. These are then trained to match the behavior of the original model. That way, we can compare the accuracy and predictions of the surrogate model to the original model. The output of each surrogate path should match the output of the convolutional downsampling it replaces under the L1 loss. The bilinear surrogate model is trained on the training data of the original model. We show that the accuracy and predictions of the surrogate model closely match those of the original model, while the generated saliency maps have less noise, as measured by total variation. And we test the saliency maps by insertion and deletion metrics introduced in [12], and observe that the bilinear surrogate model still has comparable insertion and deletion metrics compared to the original model.
@@ 2 Related Work
Saliency maps are used to highlight structures in the input that the model utilized for its prediction. The saliency methods could be divided into three groups. First, those that only utilize activations like CAM [10], secondly those that also utilize gradients of backpropagation like Grad-CAM [16], DeepLift [17], Integrated Gradients [20], or simply using the gradient [18]. And lastly pertubation base methods like RISE [12], which are usually applied at the input. Pertubation-based methods generally require several evaluations of the model, which makes them slow. To get gradient-based methods more stable SmoothGrad [19] can be employed, which runs the saliency method several times each time adding Gaussian noise to the input and then returning the average saliency map.
Gradient-based saliency methods are usually applied at either the input layer or the last convolutional layer, like [16], whereas we focus on computing the saliency maps from hidden layers of the model. [9] also computes saliency maps at hidden layers and couples them with feature visualization. To mitigate noise in the saliency maps, they employ Gaussian blurring. [14] also used Gaussian blurring to smooth saliency maps of hidden layers. We limit ourselves to models that utilize convolutional downsampling and, instead of using Gaussian blurring, we work on removing the source of the noise.
Saliency methods can be invariant of what the model has learned, which is undesired [1]. In order to measure the faithfulness of the saliency maps generated with our methods, we use deletion and insertion metrics as introduced in [12]. Using our methods, we also observe a drop in the metrics, if the model weights are randomized, suggesting the methods are sensitive to what the model has learned.
[8] visualized neurons of a CNN, and observed that the visualizations become more high-level the deeper the neuron is located in the model. We observe something similar in the attribution for the Camelyon16 and Digipath models, where in earlier layers the cell nuclei are outlined and in deeper layers segmented (Camelyon16). And the saliency map for the Digipath model highlights individual cell nuclei in earlier layers and structures formed by them in deeper layers.
@@ 2.1 Checkerboard Noise from Stride 2 Convolution
With a stride 2 convolution, not every pixel is multiplied by every weight. Due to the stride of 2, every other pixel is skipped. As an example, let $x$ be a matrix of ones with a size of $16\times 16$, $k=\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ and * the operator for a stride 2 convolution. Fig. 1 shows the result of calculating $\nabla_{x}(\sum x*k)$, i.e. the gradient towards $x$ when computing the sum of the convolution of $x$ with kernel $k$. We can see a periodic checkerboard pattern in the gradient. The black pixels were multiplied by $-1.0$ from the kernel and, therefore, received a gradient of -1.0. In contrast, the white pixels were multiplied by the $1.0$ value from the kernel and received a gradient of 1.0. That is the noise pattern we mitigate with our methods.
@@ 3.1 Overview
  • Bilinear Surrogate: Replacing each convolutional downsampling with a bilinear surrogate path (consisting of two stride 1 convolutions with a bilinear downsampling in between them), as shown in Fig. 2. This approach requires training data, in order to train the surrogate paths.
@@
In training of the bilinear surrogate model, we replace each convolutional downsampling individually (i.e. we only use the convolutional downsampling in the forward pass and not the bilinear surrogate one). But in the evaluation, we use them all together, as visible from Fig. 3. We use the L1 loss as a loss function in training.
@@ 3.2 Evaluation
With our bilinear surrogate method, we aim to achieve three goals:
@@
  • Have faithful saliency maps. If the saliency maps for the bilinear surrogate model loose faithfulness compared to the original model they would be less useful.
To validate whether we succeeded in reducing the checkerboard noise in the gradient, we can visually inspect some saliency maps. Further, we also run total variation (TV) on the saliency maps over a validation dataset. To evaluate how well we match the prediction behavior of the original model, we compare the accuracy and/or difference in prediction for only the target class and all classes of the original and the surrogate model on a validation dataset. For the difference in prediction, we use the output of the model after softmax. Then we compute the difference in prediction twice: Once only for the target class and once for all classes. To measure the faithfulness of the saliency methods we use insertion and deletion metrics introduced in [12].
@@ 3.3 Computing the Saliency Map
We use and compare four methods to compute the saliency maps, namely DeepLift, Integrated Gradients (IG), simply computing the gradient and GradCAM. We use the first three methods at both the input layer and at hidden layers, whereas we use GradCAM only at the last convolutional layer. As baseline we always use a black image (zeros), after normalization (which would correspond to a gray image for the ImageNet models before normalization). The saliency map is computed using the logits output of the model (before softmax or sigmoid).
@@ 3.4 Total Variation
$\epsilon$ is set to $10^{-6}$ and is used to avoid division by zero (i.e. if $z_{c}^{\prime}$ is zero everywhere). Eq. 2 is the resulting total variation reported in the experiments section.
@@ 4.1.1 Hardware and Software
For all our experiments, we use a Linux server with 8 Nvidia RTX 2080 Ti GPUs. On the software side, we use PyTorch [11] version 1.13.1 with CUDA version 11.6 and torchvision version 0.14.1. For computing the saliency maps we use the Captum library [6] version 0.6.0.
@@ 4.1.2 Training
The forward and backward hook approaches work on the original model, and require no additional training. All networks in our experiments are running in eval mode. We only train the bilinear surrogate paths, each consisting of two 3x3 stride 1 convolutions, with a bilinear downsampling in between them. The rest of the network is kept frozen. We train the bilinear surrogate for 10 epochs. As optimizer we use Adam [5] with a learning rate of 0.001. As batch size we use 64 for training the surrogate models on the ImageNet1K [2] data, 512 for the Camelyon16 data and 16 for the Digipath data (due to the larger patch size of 1536 we use a smaller batch size here).
@@
Training Samples 1 281 16712811671\,281\,1671 281 167 9 655 75696557569\,655\,7569 655 756 129 634129634129\,634129 634
Validation Samples 50 0005000050\,00050 000 46 267 8994626789946\,267\,89946 267 899 $899$
@@ 4.1.3 Datasets
  • ImageNet1K. We evaluate the accuracy and the difference in prediction between the original and surrogate model on the 50 0005000050\,00050 000 validation images. The total variation and saliency metrics are evaluated on every hundredth image of the validation data, resulting in 500 images.
  • Camelyon16. For training we use all the patches of the original training data were the original model predicts tumor (score after sigmoid > 0), plus randomly sample the same amount of patches were the original model is not predicting tumor (score after sigmoid < 0). The difference in prediction is evaluated on all the 46 267 8994626789946\,267\,89946 267 899 validation samples. The total variation and saliency metrics are evaluated on 500 samples were the original model predicts tumor randomly chosen out of the validation samples.
@@ 4.1.4 Models
On the ImageNet1K data we use a ResNet34 and a ResNet50 both openly available from the pytorch model zoo. Further, we use an adversarially robust ResNet50 from [15]. For the Camelyon16 data we use an openly available ResNet18 from the monai model zoo. And for the Digipath data we use an in-house U-Net segmentation model with a ResNet34 backbone. While introduced a while ago, the ResNet50 architecture is still utilized in recent work [13].
4.2 Accuracy and Prediction Difference
@@
Tab. 2 shows the accuracy of the original, bilinear surrogate and forward hooked models, on the ImageNet1K data. For the bilinear surrogate models we observe a maximum decrease of accuracy of 0.17 for the ResNet34 model, and a minimum decrease of 0.03 for the robust ResNet50 model. The forward hooked models have a larger decrease of accuracy, especially in the case of the ResNet34 model where the accuracy decreases by 20.04.
@@
In Tab. 3 we evaluate how the prediction of the models changes, once for the sum over all predicted classes and once only for the target class. We report the differences in softmax values times 100 (i.e. a reported difference of 9.52 for the ResNet34 model would mean a difference of 0.0952). For the bilinear surrogate models we observe the maximum change of prediction for the sum of all classes and only for the target class of 9.52 respectively 2.39 for the Resnet34 model. That aligns with the drop of accuracy being largest for the ResNet34 surrogate model. Averaged over all models, the forward hooked models have a 7.8 times higher difference in prediction for all classes, and a 8.8 times higher difference in prediction for the target class compared to the bilinear surrogate model.
Table 4: Best insertion and deletion layers and methods.
Network Score DeepLift IG Grad GradCAM
ResNet34 Ins $\uparrow$ 0.679 (S)(2_3) $0.663$ (S)(2_3) $0.559$ (F)(4_0) $0.648$
Del $\downarrow$ 0.045 (O)(1_2) $0.052$ (O)(1_2) $0.081$ (O)(1_2) $0.127$
ResNet50 Ins 0.706 (S)(3_1) $0.683$ (S)(4_2) $0.531$ (B)(3_4) $0.684$
Del 0.044 (O)(1_2) $0.052$ (O)(1_2) $0.071$ (O)(1_1) $0.147$
ResNet50 Robust Ins $0.492$ (O)(Input) 0.544 (O)(Input) $0.430$ (O)(Input) $0.397$
Del 0.041 (O)(1_1) $0.042$ (O)(Input) $0.070$ (O)(3_2) $0.066$
ResNet18 Camelyon Ins $0.738$ (S)(1_1) 0.775 (O)(1_0) 0.630 (O)(1_2) $0.625$
Del 0.022 (O)(Input) $0.030$ (O)(Input) 0.062 (O)(2_0) $0.366$
(a) ResNet34 del $\downarrow$
(b) ResNet50 del $\downarrow$
(c) ResNet18 del $\downarrow$
(d) ResNet34 ins $\uparrow$
(e) ResNet50 ins $\uparrow$
(f) ResNet18 ins $\uparrow$
Figure 4: Insertion and Deletion metrics using DeepLift.
4.3 Insertion and Deletion
In this section we evaluate the saliency maps based on the insertion and deletion metrics proposed in [12] and use their implementation. That is, for deletion we set the deleted pixels to gray ((0, 0, 0) after normalization) for the ImageNet and Camelyon16 models (same value that we used as baseline in the attribution methods). For insertion we use a blurred version of the original image as starting point.
Fig. 4 shows the insertion and deletion scores for DeepLift for the ResNet34, ResNet50 and ResNet18 Camelyon16 models. On the x-axis is the layer, going deeper into the model from left to right. As baselines we use DeepLift computed towards the input layer, and GradCAM computed at the last convolutional layer (layer 4_2) both times using the original model (results for the other saliency methods as well as for the robust ResNet50 can be found in the supplementary material).
Tab. 4 shows the layer and method that performs best for insertion and deletion for each of the four saliency methods. Each entry can be read as: Metric result (method)(layer). The method can be one of four: Original model (O), bilinear surrogate (S), backward hooked (B), forward hooked (F). For GradCAM we always use the original model and the last convolutional layer. In total either DeepLift or Integrated Gradient performs best, with DeepLift performing best 6 out of 8 times. Further attribution in hidden layers performs better than in the input layer (6 out of 8 times). Of those 6 times where a hidden layer performs best, twice the bilinear surrogate model performs best (both times in insertion).
Averaged over the ImageNet1K models and layers, the attribution with DeepLift in hidden layers for the bilinear surrogate model has a 2.14% higher insertion score compared to the original model. But it also has an 11.8% higher deletion score compared to the original model. The reason for the worse deletion score could likely be that in the original model some pixels count more than others due to the convolutional downsampling. Therefore the model may be relying on less input pixels for making its predictions, whereas the bilinear surrogate relies on the pixels more evenly. The backward hook method has an 0.84% higher insertion score, but a 16.2% higher deletion score. And the forward hook method has both a 7.02% lower insertion score, and a 44.2% higher deletion score, making it considerably less faithful compared to using the original model.
4.4 Randomize Model Weights
The result of saliency methods should depend on what the model has learned, since we want to explain the trained model. [1] found that some saliency methods, are invariant of what the model has learned and rather depend on the architecture, those saliency methods give similar results for trained and randomly initialized models. Therefore, we test our methods with the model randomization test. We set the layers of the trained model to the random initialized version, starting from the end of the model (from the fully connected layer).
Deletion Layer1
Insertion Layer1
Figure 5: Insertion and Deletion scores for randomizing the model, using DeepLift as attribution method.
Fig. 5 shows the results for the ResNet34 model on ImageNet1K data. The figure shows the insertion and deletion scores when computing the attribution at Layer1 using DeepLift. The bottom of each sub-figure shows till which layer the model is randomized. The metrics generally become worse, the more layers are randomized. For the deletion metrics, using a completely randomized model still performs a bit better compared to using a random saliency map (gaussian noise at the spatial size of the layer, upscaled to the input size), suggesting the model architecture has some impact on the result. But overall the metrics become considerably worse when the model is randomized, which we see as confirmation that the method is sensitive to what the model has learned.